OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6? Explain.

OpenStudy (ranga):

Put the equation of the line, 3x – 2y = –6, in slope-intercept form and find the slope of the line. Parallel line will have the same slope. The equation of a line passing through (4,2) is: (y-2) = m(x-4). plug m with slope found earlier.

OpenStudy (the_fizicx99):

3x – 2y = –6 -3x -3x -2y = -3x - 6 /-2 /-2 /-2 y = 3/2x +3 M = 3/2 Now plug (4,2) to, \[y-y1=m(x-x1)\]